# Tree metrics and log-concavity for matroids

### v1

@@ Abstract.
We show that a set function $\nu$ satisfies the gross substitutes property if and only if its homogeneous generating polynomial $Z_{q,\nu}$ is a Lorentzian polynomial for all positive $q\leq 1$, answering a question of Eur–Huh. We achieve this by giving a rank $1$ upper bound for the distance matrix of an ultrametric tree, refining a classical result of Graham–Pollak. This characterization enables us to resolve two open problems that strengthen Mason’s log-concavity conjectures for the numbers of independent sets of a matroid: one posed by Giansiracusa–Rincón–Schleis–Ulirsch for valuated matroids, and another posed by Pak for ordinary matroids.
@@ 1. Introduction
The rank of $\mathsf{M}$ is the common cardinality of the maximal independent sets. For background and any undefined terminology in matroid theory, we refer to [18].
Let $I_{k}=I_{k}(\mathsf{M})$ be the number of independent sets of $\mathsf{M}$ of cardinality $k$. Mason [14] conjectured that the following three families of inequalities hold for all $0<k<n$:
@@
Conjecture (M1) was proved for realizable matroids in [12], by building on the work of [9], and for arbitrary matroids in [1]. Conjecture (M2) was proved in [10] using the main results of [1], while Conjecture (M3) was proved independently in [2] and [4]. These inequalities motivated the development of combinatorial Hodge theory [1] and the theory of Lorentzian polynomials [4].
In Theorem 1.4, we extend (M3) from matroids on $E$ to $\text{M}^{\natural}$-concave functions on $2^{E}$, which are precisely the set functions satisfying the gross substitutes property from economics. This answers a question of Giansiracusa–Rincón–Schleis–Ulirsch in [7, Question 4.1], who asked whether (M3) holds for the class of $\text{M}^{\natural}$-concave functions constructed from valuated matroids discussed in Example 1.3. In Theorem 1.6, we prove a multivariate polynomial inequality that refines (M2), affirmatively answering a question posed by Pak [19]. Both results follow from our main result, Theorem 1.5, characterizing $\text{M}^{\natural}$-concave functions in terms of Lorentzian polynomials:
@@
This answers a question of Eur–Huh in [5, Section 5.1], who asked for a characterization of set functions whose homogeneous generating polynomials are Lorentzian for all positive $q\leq 1$. [^1] The function $r$ from [5, Example 5.4] is not $\text{M}^{\natural}$-concave, contrary to the authors’ claim: the maximum of $r(\left\{i,j\right\})+r(\left\{k\right\})$ is achieved uniquely at $k=2$. The key ingredient is Theorem 1.8, which gives a rank $1$ upper bound for the distance matrix of an ultrametric tree. This refines the classical result of Graham–Pollak that the distance matrix of a tree has exactly one positive eigenvalue. In the rest of the introduction, we describe these results in detail.
@@ 1.1. Inequalities for $\text{M}^{\natural}$-concave set functions
Introduced by Murota and Shioura [17], $\text{M}^{\natural}$-concave functions are central objects in discrete convex analysis. Fujishige–Yang [6] and Reijnierse–van Gellekom–Potters [20] independently proved that $\text{M}^{\natural}$-concavity is equivalent to the gross substitutes property in economics, introduced by Kelso and Crawford [11] two decades earlier. For a comprehensive introduction to $\text{M}^{\natural}$-concave functions, we refer the reader to [16, Chapter 6]. The notion of $\text{M}^{\natural}$-concave functions generalizes several concepts in matroid theory, including independent sets, rank functions, and valuated matroids.
@@
is an $\text{M}^{\natural}$-concave function whose effective domain is $2^{E}$. More generally, any non-negative linear combination of the rank functions of the constituent matroids in a flag matroid is a $\text{M}^{\natural}$-concave function [22, Theorem 3].
@@
Valuated matroids are precisely the possible height functions in a regular subdivision of a matroid polytope into matroid polytopes. [23]
Murota [15] considered the $\text{M}^{\natural}$-concave extension $\nu$ of $\underline{\nu}$ to $2^{E}$ given by
@@
$I_{q,\nu;k}\coloneq\sum_{I\in\binom{E}{k}}q^{-\nu(I)},$
where, by convention, $q^{\infty}=0$. This sequence was first considered by Giansiracusa, Rincón, Schleis, and Ulirsch for valuated matroids in the context of Example 1.3, who proved in [7, Theorem A] that Murota’s extension $\nu$ of a valuated matroid satisfies a generalization of (M2):
@@
They asked in [7, Question 4.1] whether $\nu$ satisfies a generalization of (M3), and they provided extensive numerical evidence in support of this. We prove their prediction in the more general setting of $\text{M}^{\natural}$-concave functions.
@@
We deduce Theorem 1.4 from an analytic characterization of $\text{M}^{\natural}$-concave functions in terms of Lorentzian polynomials [4], whose definition we recall in Section 3.1. We define the homogeneous generating polynomial a function $\nu\colon 2^{E}\to\mathbb{R}\cup\{-\infty\}$ by
$Z_{q,\nu}(\mathsf{x},y)\coloneq\sum_{S\subseteq E}q^{-\nu(S)}\mathsf{x}^{S}y^{\left|E\setminus S\right|}\ \ \text{with}\ \ \mathsf{x}^{S}\coloneq\prod_{i\in S}x_{i},$
@@
Theorem 1.5 answers a question of Eur–Huh in [5, Section 5], who proved the “if” direction of Theorem 1.5 in [5, Proposition 5.5]. The other direction is a new contribution, which generalizes the following known results:
@@
    $I(\mathsf{x},y)\coloneq\sum\limits_{I\in\operatorname{{IN}}(\mathsf{M})}\mathsf{x}^{I}y^{\left|E\setminus I\right|},$
    is a Lorentzian polynomial [4, Section 4.3]. See [2, Theorem 4.1] for an equivalent statement.
@@
    $T_{q}(\mathsf{x},y)\coloneq\sum\limits_{S\subseteq E}q^{-\operatorname{rk}_{\mathsf{M}}(S)}\mathsf{x}^{S}y^{\left|E\setminus S\right|},$
    is a Lorentzian polynomial for all positive $q\leq 1$ in [4, Theorem 4.10]. The Lorentzian property of the homogeneous independent set generating polynomial follows from the identity
@@ 1.2. Polynomial inequalities for matroids
$I_{k}(\mathsf{x})\coloneqq\sum_{S\in\operatorname{{IN}}(\mathsf{M})_{k}}\mathsf{x}^{S},\ \ \text{with}\ \ \mathsf{x}^{S}\coloneq\prod_{i\in S}x_{i},$
where $\mathsf{x}=(x_{i})_{i\in E}$. For multivariate polynomials $f,g\in\mathbb{Z}[x_{i}]_{i\in E}$, we write $f\succeq g$ if all the coefficients of the difference fgf-g are nonnegative. Pak [19] asked whether the polynomial refinement of Mason’s conjecture (M1) holds for any matroid: Is the inequality $I_{k}(\mathsf{x})^{2}\succeq I_{k-1}(\mathsf{x})I_{k+1}(\mathsf{x})$ true for all $k$, coefficient by coefficient?
We answer Pak’s question affirmatively. Using Theorem 1.5, we prove a stronger polynomial inequality for the more general class of $\text{M}^{\natural}$-concave functions. For a function $\nu:2^{E}\to\mathbb{R}\cup\{-\infty\}$, we set
$I_{q,\nu;k}(\mathsf{x})\coloneq\sum_{S\in\binom{E}{k}}q^{-\nu(S)}\mathsf{x}^{S}.$
When $\nu$ is the independent set indicator function of $\mathsf{M}$, the following theorem states that the polynomial version of (M2) holds for $\mathsf{M}$. This answers Pak’s question on (M1).
@@ 1.3. An inequality for ultrametric trees
A central ingredient in the proof of Theorem 1.5 is a result on tree distance matrices that we now describe. In their design of efficient address systems for communication networks, Graham and Pollak [8] introduced the distance matrix of a graph, whose $ij$ entry is the distance between vertices $i$ and $j$. They showed that the signature of this matrix gives a lower bound for the addresses in their design, and proved that the distance matrix of any tree has the Lorentzian signature $(+,-,\cdots,-)$. The latter fact also follows from work of Schoenberg [21, Theorem 1], using the fact that an ultrametric tree can be metrically embedded in an $\ell^{2}$-space [24]. Our proofs of the matroid inequalities above rely on a refinement of this result for ultrametric trees.
@@
Theorem 1.8 serves as the base case in our inductive argument for the proof of Theorem 1.5. This connection stems from Lemma 3.3, which states that any $\text{M}^{\natural}$-concave function on $2^{E}$ whose effective domain contains all the subsets of $E$ with at most $2$ elements gives rise to an ultrametric tree. This is a variation of the standard fact that the space of uniform valuated matroids of rank 2 is equal to the space of phylogenetic trees [13, Theorem 4.3.5].
@@ Acknowledgements
The authors gratefully acknowledge the Institute for Advanced Study for providing an inspiring research environment. Federico Ardila–Mantilla was supported by NSF Grant DMS 2154279 and a Friends of the Institute for Advanced Study Membership. Graham Denham was supported by NSERC of Canada. Chris Eur was supported by NSF Grant DMS 2246518. Sergio Cristancho and June Huh were supported by the Simons Investigator Grant. Botong Wang was supported by NSF Grant DMS 1926686.
@@ 2.1. Ultrametric tree matrices
Every ultrametric tree with positive lengths defines an ultrametric on its leaves, and every finite ultrametric space arises in this way from an ultrametric tree [3, Theorem 3.1].
@@
(2.1) $i\vee k=j\vee k=h\vee k\ \ \text{for any $k\in M\backslash\left\{i,j\right\}$}.$
Setting $\langle a\rangle\coloneq 1-\frac{1}{a}$ and focusing on the $i$-th and $j$-th rows and columns of the matrix $A^{T,U}$, which are shown as the first and the second rows and columns below, we have
@@
[…]
$\sim$ $\begin{bmatrix}&(2-\left<n_{i}\right>-\left<n_{j}\right>)H(h)&0&0&\cdots\\ &0&\left<n\right>-\left<n_{h}\right>H(h)&\left<n\right>-H(h\vee k)&\cdots\\ &0&\left<n\right>-H(h\vee k)&\left<n\right>-\left<n_{k}\right>H(k)&\cdots\\ &\vdots&\vdots&\vdots&\ddots&\end{bmatrix}{\begin{matrix}\text{replacing}\\ \text{row }j&\mapsto&\text{row }j+\frac{n_{i}}{n_{h}}\text{row }i\\ \text{col }j&\mapsto&\text{col }j+\frac{n_{i}}{n_{h}}\text{col }i\end{matrix}}$
@@ 3.1. Lorentzian polynomials and discrete convex analysis
$\partial^{\alpha}\coloneq\prod_{i\in E}\partial_{i}^{\alpha_{i}}\ \ \text{for any $\alpha=(\alpha_{i})_{i\in E}\in\mathbb{Z}^{E}_{\geq 0}$}.$
@@
$\mathring{\mathrm{L}}^{2}_{n}\coloneq\left\{f\in\mathrm{H}^{2}_{n}\;\middle|\;\left(\partial_{i}\partial_{j}f\right)_{i,j\in E}\text{~has only positive entries and signature}(+,-\cdots,-)\right\}.$
@@
$\mathring{\mathrm{L}}^{d}_{n}\coloneq\left\{f\in\mathrm{H}^{d}_{n}\;\middle|\;\partial_{i}f\in\mathring{\mathrm{L}}_{n}^{d-1}\text{~for all $i\in E$}\right\}.$
@@
    If $f(\mathsf{x})\in\mathrm{L}^{d}_{n}$ and $A$ is an $n\times m$ nonnegative matrix, then $f(A\mathsf{y})\in\mathrm{L}^{d}_{m}$ [4, Theorem 2.10].
@@
    If $f\in\mathrm{L}_{n}^{d}$, then $\sum_{i}a_{i}\partial_{i}f\in\mathrm{L}_{n}^{d-1}$ for any $a_{i}\geq 0$ [4, Corollary 2.11].
@@
    If $f\in\mathrm{L}_{n}^{d}$ and $g\in\mathrm{L}_{n}^{e}$, then $fg\in\mathrm{L}^{d+e}_{n}$ [4, Corollary 2.32].
@@
    is a log-concave sequence of nonnegative numbers with no internal zeros [4, Example 2.3].
We briefly review the theory of Lorentzian polynomials and its relation to discrete convex analysis. For further details, we refer to [4] for Lorentzian polynomials and [16] for discrete convex analysis. We consider the rank $d$ simplex in $\mathbb{Z}_{\geq 0}^{E}$ defined by
@@
$\operatorname{{dom}}(\nu)\coloneq\left\{\alpha\in\Delta_{n}^{d}\;\middle|\;\nu(\alpha)\neq\infty\right\},$
@@
$\operatorname{supp}(f)\coloneq\left\{\alpha\in\Delta_{n}^{d}\;\middle|\;c_{\alpha}\neq 0\right\}.$
The following characterization of Lorentzian polynomials is central to their relationship to matroids and M-convexity [4, Theorem 2.25].
@@
From this, one may deduce the following characterizations of M-convex functions [4, Theorem 3.14]: A function $\nu:\Delta^{d}_{n}\to\mathbb{R}\cup\{\infty\}$ is M-convex if and only if its normalized generating polynomial
$f_{q,\nu}(\mathsf{x})\coloneq\sum_{\alpha\in\operatorname{{dom}}(\nu)}q^{\nu(\alpha)}\frac{\mathsf{x}^{\alpha}}{\alpha!}$
is a Lorentzian polynomial for all positive $q\leq 1$, where ${\mathsf{x}^{\alpha}}/{\alpha!}$ is the product of ${x_{i}^{\alpha_{i}}}/{\alpha_{i}!}$ for $i\in E$. This in turn implies the following characterization of M-convex sets [4, Theorem 3.10]: A subset $\mathrm{J}$ of $\Delta^{d}_{n}$ is M-convex if and only if its normalized generating polynomial
$f_{\mathrm{J}}\coloneq\sum_{\alpha\in\mathrm{J}}\frac{\mathsf{x}^{\alpha}}{\alpha!}$
@@
$N(Z_{q,\nu})\coloneq\sum_{S\subseteq E}q^{-\nu(S)}\mathsf{x}^{S}\frac{y^{\left|E\setminus S\right|}}{\left|E\setminus S\right|!}$

### v2

@@ Abstract.
We show that a set function $\nu$ satisfies the gross substitutes property if and only if its homogeneous generating polynomial $Z_{q,\nu}$ is a Lorentzian polynomial for all positive $q\leq 1$, answering a question of Eur–Huh. We achieve this by giving a rank $1$ upper bound for the distance matrix of an ultrametric tree, refining a classical result of Graham–Pollak. This characterization enables us to resolve open problems that strengthen Mason’s log-concavity conjectures for the numbers of independent sets of a matroid: one posed by Giansiracusa–Rincón–Schleis–Ulirsch for valuated matroids, and two posed by Dowling in 1980 and Zhao in 1985 for ordinary matroids.
@@ 1. Introduction
The rank of $\mathsf{M}$ is the common cardinality of the maximal independent sets. For background and any undefined terminology in matroid theory, we refer to [Oxl11].
Let $I_{k}=I_{k}(\mathsf{M})$ be the number of independent sets of $\mathsf{M}$ of cardinality $k$. Mason [Mas72] conjectured that the following three families of inequalities hold for all $0<k<n$:
@@
Conjecture (M1) was proved for realizable matroids in [Len13], by building on the work of [HK12], and for arbitrary matroids in [AHK18]. Conjecture (M2) was proved in [HSW22] using the main results of [AHK18], while Conjecture (M3) was proved independently in [ALOGV24] and [BH20]. These inequalities motivated the development of combinatorial Hodge theory [AHK18] and the theory of Lorentzian polynomials [BH20].
In Theorem 1.4, we extend (M3) from matroids on $E$ to $\text{M}^{\natural}$-concave functions on $2^{E}$, which are precisely the set functions satisfying the gross substitutes property from economics. This answers a question of Giansiracusa–Rincón–Schleis–Ulirsch in [GRSU24, Question 4.1], who asked whether (M3) holds for the class of $\text{M}^{\natural}$-concave functions constructed from valuated matroids discussed in Example 1.3. In Theorem 1.6, we prove a multivariate polynomial inequality that refines (M2), affirmatively settling conjectures of Dowling [Dow80] and Zhao [Zha85] that were recently popularized by Pak [Pak24]. Both results follow from our main result, Theorem 1.5, characterizing $\text{M}^{\natural}$-concave functions in terms of Lorentzian polynomials:
@@
This answers a question of Eur–Huh in [EH20, Section 5.1], who asked for a characterization of set functions whose homogeneous generating polynomials are Lorentzian for all positive $q\leq 1$. [^1] The function $r$ from [EH20, Example 5.4] is not $\text{M}^{\natural}$-concave, contrary to the authors’ claim: the maximum of $r(\left\{i,j\right\})+r(\left\{k\right\})$ is achieved uniquely at $k=2$. The key ingredient is Theorem 1.8, which gives a rank $1$ upper bound for the distance matrix of an ultrametric tree. This refines the classical result of Graham–Pollak that the distance matrix of a tree has exactly one positive eigenvalue. In the rest of the introduction, we describe these results in detail.
@@ 1.1. Inequalities for $\text{M}^{\natural}$-concave set functions
Introduced by Murota and Shioura [MS99], $\text{M}^{\natural}$-concave functions are central objects in discrete convex analysis. Fujishige–Yang [FY03] and Reijnierse–van Gellekom–Potters [RvGP02] independently proved that $\text{M}^{\natural}$-concavity is equivalent to the gross substitutes property in economics, introduced by Kelso and Crawford [KJC82] two decades earlier. For a comprehensive introduction to $\text{M}^{\natural}$-concave functions, we refer the reader to [Mur03, Chapter 6]. The notion of $\text{M}^{\natural}$-concave functions generalizes several concepts in matroid theory, including independent sets, rank functions, and valuated matroids.
@@
is an $\text{M}^{\natural}$-concave function whose effective domain is $2^{E}$. More generally, any non-negative linear combination of the rank functions of the constituent matroids in a flag matroid is a $\text{M}^{\natural}$-concave function [Shi12, Theorem 3].
@@
Valuated matroids are precisely the possible height functions in a regular subdivision of a matroid polytope into matroid polytopes. [Spe08]
Murota [Mur97] considered the $\text{M}^{\natural}$-concave extension $\nu$ of $\underline{\nu}$ to $2^{E}$ given by
@@
$I_{q,\nu;k}\coloneqq\sum_{I\in\binom{E}{k}}q^{-\nu(I)},$
where, by convention, $q^{\infty}=0$. This sequence was first considered by Giansiracusa, Rincón, Schleis, and Ulirsch for valuated matroids in the context of Example 1.3, who proved in [GRSU24, Theorem A] that Murota’s extension $\nu$ of a valuated matroid satisfies a generalization of (M2):
@@
They asked in [GRSU24, Question 4.1] whether $\nu$ satisfies a generalization of (M3), and they provided extensive numerical evidence in support of this. We prove their prediction in the more general setting of $\text{M}^{\natural}$-concave functions.
@@
We deduce Theorem 1.4 from an analytic characterization of $\text{M}^{\natural}$-concave functions in terms of Lorentzian polynomials [BH20], whose definition we recall in Section 3.1. We define the homogeneous generating polynomial a function $\nu\colon 2^{E}\to\mathbb{R}\cup\{-\infty\}$ by
$Z_{q,\nu}(\mathsf{x},y)\coloneqq\sum_{S\subseteq E}q^{-\nu(S)}\mathsf{x}^{S}y^{\left|E\setminus S\right|}\ \ \text{with}\ \ \mathsf{x}^{S}\coloneqq\prod_{i\in S}x_{i},$
@@
Theorem 1.5 answers a question of Eur–Huh in [EH20, Section 5], who proved the “if” direction of Theorem 1.5 in [EH20, Proposition 5.5]. The other direction is a new contribution, which generalizes the following known results:
@@
    $I(\mathsf{x},y)\coloneqq\sum\limits_{I\in\operatorname{{IN}}(\mathsf{M})}\mathsf{x}^{I}y^{\left|E\setminus I\right|},$
    is a Lorentzian polynomial [BH20, Section 4.3]. See [ALOGV24, Theorem 4.1] for an equivalent statement.
@@
    $T_{q}(\mathsf{x},y)\coloneqq\sum\limits_{S\subseteq E}q^{-\operatorname{rk}_{\mathsf{M}}(S)}\mathsf{x}^{S}y^{\left|E\setminus S\right|},$
    is a Lorentzian polynomial for all positive $q\leq 1$ in [BH20, Theorem 4.10]. The Lorentzian property of the homogeneous independent set generating polynomial follows from the identity
@@ 1.2. Polynomial inequalities for matroids
$I_{k}(\mathsf{x})\coloneqq\sum_{S\in\operatorname{{IN}}(\mathsf{M})_{k}}\mathsf{x}^{S},\ \ \text{with}\ \ \mathsf{x}^{S}\coloneqq\prod_{i\in S}x_{i},$
where $\mathsf{x}=(x_{i})_{i\in E}$. For multivariate polynomials $f,g\in\mathbb{Z}[x_{i}]_{i\in E}$, we write $f\succeq g$ if all the coefficients of the difference fgf-g are nonnegative. Dowling [Dow80, Section 2] conjectured that the polynomial refinement of Mason’s conjecture (M1) holds for any matroid: that is, the inequality $I_{k}(\mathsf{x})^{2}\succeq I_{k-1}(\mathsf{x})I_{k+1}(\mathsf{x})$ holds for all $k$, coefficient by coefficient. Subsequently, Zhao [Zha85] made the stronger conjecture that the polynomial version of (M2) holds for all matroids.
We prove Zhao’s and Dowling’s conjectures. Using Theorem 1.5, we establish a stronger polynomial inequality for the more general class of $\text{M}^{\natural}$-concave functions. For a function $\nu:2^{E}\to\mathbb{R}\cup\{-\infty\}$, we set
$I_{q,\nu;k}(\mathsf{x})\coloneqq\sum_{S\in\binom{E}{k}}q^{-\nu(S)}\mathsf{x}^{S}.$
When $\nu$ is the independent set indicator function of $\mathsf{M}$, the following theorem states that the polynomial version of (M2), and hence also (M1), holds for $\mathsf{M}$.
@@ 1.3. An inequality for ultrametric trees
A central ingredient in the proof of Theorem 1.5 is a result on tree distance matrices that we now describe. In their design of efficient address systems for communication networks, Graham and Pollak [GP71] introduced the distance matrix of a graph, whose $ij$ entry is the distance between vertices $i$ and $j$. They showed that the signature of this matrix gives a lower bound for the addresses in their design, and proved that the distance matrix of any tree has the Lorentzian signature $(+,-,\cdots,-)$. The latter fact also follows from work of Schoenberg [Sch35, Theorem 1], using the fact that an ultrametric tree can be metrically embedded in an $\ell^{2}$-space [TV83]. Our proofs of the matroid inequalities above rely on a refinement of this result for ultrametric trees.
@@
Theorem 1.8 serves as the base case in our inductive argument for the proof of Theorem 1.5. This connection stems from Lemma 3.3, which states that any $\text{M}^{\natural}$-concave function on $2^{E}$ whose effective domain contains all the subsets of $E$ with at most $2$ elements gives rise to an ultrametric tree. This is a variation of the standard fact that the space of uniform valuated matroids of rank 2 is equal to the space of phylogenetic trees [MS15, Theorem 4.3.5].
@@ Acknowledgements
The authors gratefully acknowledge the Institute for Advanced Study for providing an inspiring research environment. Federico Ardila–Mantilla was supported by NSF Grant DMS 2154279, the Friends of the Institute for Advanced Study Fund, and a UK Royal Society Wolfson Fellowship. Graham Denham was supported by NSERC of Canada. Chris Eur was supported by NSF Grant DMS 2246518. Sergio Cristancho and June Huh were supported by the Simons Investigator Grant. Botong Wang was supported by NSF Grant DMS 1926686.
Our results and proofs were first announced in Feb. 2025 at the Institute for Advanced Study [Ard25a, Ard25b]. In Jan. 2026, a few days after our work appeared on the arXiv, Cao, Chen, Li, and Wu [CCLW26] announced results that overlap with one of our four main theorems. We thank them for bringing Dowling and Zhao’s conjectures to our attention.
@@ 2.1. Ultrametric tree matrices
Every ultrametric tree with positive lengths defines an ultrametric on its leaves, and every finite ultrametric space arises in this way from an ultrametric tree [BG91, Theorem 3.1].
@@
$i\vee k=j\vee k=h\vee k\ \ \text{for any $k\in M\backslash\left\{i,j\right\}$}.$
Setting $\langle a\rangle\coloneqq 1-\frac{1}{a}$ and focusing on the $i$-th and $j$-th rows and columns of the matrix $A^{T,U}$, which are shown as the first and the second rows and columns below, we have
@@
$\sim$ $\begin{bmatrix}&(2-\left<n_{i}\right>-\left<n_{j}\right>)H(h)&(\left<n_{j}\right>-1)H(h)&0&\cdots\\ &(\left<n_{j}\right>-1)H(h)&\left<n\right>-\left<n_{j}\right>H(h)&\left<n\right>-H(h\vee k)&\cdots\\ &0&\left<n\right>-H(h\vee k)&\left<n\right>-\left<n_{k}\right>H(k)&\cdots\\ &\vdots&\vdots&\vdots&\ddots&\end{bmatrix}{\begin{aligned} &\text{replacing}\\ &\text{row }i\mapsto\text{row }i-\text{row }j\\ &\text{col }i\mapsto\text{col }i-\text{col }j,\\ &\text{using $i\vee k=j\vee k=h\vee k$.}\end{aligned}}$
$\sim$ $\begin{bmatrix}&(2-\left<n_{i}\right>-\left<n_{j}\right>)H(h)&0&0&\cdots\\ &0&\left<n\right>-\left<n_{h}\right>H(h)&\left<n\right>-H(h\vee k)&\cdots\\ &0&\left<n\right>-H(h\vee k)&\left<n\right>-\left<n_{k}\right>H(k)&\cdots\\ &\vdots&\vdots&\vdots&\ddots&\end{bmatrix}{\begin{aligned} &\text{replacing}\\ &\text{row }j\mapsto\text{row }j+\frac{n_{i}}{n_{h}}\text{row }i\\ &\text{col }j\mapsto\text{col }j+\frac{n_{i}}{n_{h}}\text{col }i.\end{aligned}}$
@@ 3.1. Lorentzian polynomials and discrete convex analysis
$\partial^{\alpha}\coloneqq\prod_{i\in E}\partial_{i}^{\alpha_{i}}\ \ \text{for any $\alpha=(\alpha_{i})_{i\in E}\in\mathbb{Z}^{E}_{\geq 0}$}.$
@@
$\mathring{\mathrm{L}}^{2}_{n}\coloneqq\left\{f\in\mathrm{H}^{2}_{n}\;\middle|\;\left(\partial_{i}\partial_{j}f\right)_{i,j\in E}\text{~has only positive entries and signature}(+,-\cdots,-)\right\}.$
@@
$\mathring{\mathrm{L}}^{d}_{n}\coloneqq\left\{f\in\mathrm{H}^{d}_{n}\;\middle|\;\partial_{i}f\in\mathring{\mathrm{L}}_{n}^{d-1}\text{~for all $i\in E$}\right\}.$
@@
    If $f(\mathsf{x})\in\mathrm{L}^{d}_{n}$ and $A$ is an $n\times m$ nonnegative matrix, then $f(A\mathsf{y})\in\mathrm{L}^{d}_{m}$ [BH20, Theorem 2.10].
@@
    If $f\in\mathrm{L}_{n}^{d}$, then $\sum_{i}a_{i}\partial_{i}f\in\mathrm{L}_{n}^{d-1}$ for any $a_{i}\geq 0$ [BH20, Corollary 2.11].
@@
    If $f\in\mathrm{L}_{n}^{d}$ and $g\in\mathrm{L}_{n}^{e}$, then $fg\in\mathrm{L}^{d+e}_{n}$ [BH20, Corollary 2.32].
@@
    is a log-concave sequence of nonnegative numbers with no internal zeros [BH20, Example 2.3].
We briefly review the theory of Lorentzian polynomials and its relation to discrete convex analysis. For further details, we refer to [BH20] for Lorentzian polynomials and [Mur03] for discrete convex analysis. We consider the rank $d$ simplex in $\mathbb{Z}_{\geq 0}^{E}$ defined by
@@
$\operatorname{{dom}}(\nu)\coloneqq\left\{\alpha\in\Delta_{n}^{d}\;\middle|\;\nu(\alpha)\neq\infty\right\},$
@@
$\operatorname{supp}(f)\coloneqq\left\{\alpha\in\Delta_{n}^{d}\;\middle|\;c_{\alpha}\neq 0\right\}.$
The following characterization of Lorentzian polynomials is central to their relationship to matroids and M-convexity [BH20, Theorem 2.25].
@@
From this, one may deduce the following characterizations of M-convex functions [BH20, Theorem 3.14]: A function $\nu:\Delta^{d}_{n}\to\mathbb{R}\cup\{\infty\}$ is M-convex if and only if its normalized generating polynomial
$f_{q,\nu}(\mathsf{x})\coloneqq\sum_{\alpha\in\operatorname{{dom}}(\nu)}q^{\nu(\alpha)}\frac{\mathsf{x}^{\alpha}}{\alpha!}$
is a Lorentzian polynomial for all positive $q\leq 1$, where ${\mathsf{x}^{\alpha}}/{\alpha!}$ is the product of ${x_{i}^{\alpha_{i}}}/{\alpha_{i}!}$ for $i\in E$. This in turn implies the following characterization of M-convex sets [BH20, Theorem 3.10]: A subset $\mathrm{J}$ of $\Delta^{d}_{n}$ is M-convex if and only if its normalized generating polynomial
$f_{\mathrm{J}}\coloneqq\sum_{\alpha\in\mathrm{J}}\frac{\mathsf{x}^{\alpha}}{\alpha!}$
@@
$N(Z_{q,\nu})\coloneqq\sum_{S\subseteq E}q^{-\nu(S)}\mathsf{x}^{S}\frac{y^{\left|E\setminus S\right|}}{\left|E\setminus S\right|!}$